OpenStudy (anonymous):

A 5ft tall person is walking away from a 17ft tall lamppost at a rate of 6ft/sec. The scenario can be modeled with right triangles. At what rate is the length of the person's shadow changing when the person is 16ft from the lamppost?

OpenStudy (anonymous):



OpenStudy (anonymous):

you can use \[\frac{17}{5}=\frac{x+y}{y}\]

OpenStudy (anonymous):

this gives you a nice linear relation between \(x\) and \(y\) namely \[12y=5x\]

OpenStudy (anonymous):

since it is linear, the 16 part is unimportant since you have \[y=\frac{5}{12}x\] you know \[y'=\frac{5}{12}x'\] and you also know \(x'=6\) so you get \[y'=\frac{30}{12}=\frac{5}{2}=2.5\]

OpenStudy (anonymous):

GAH. I drew my triangles wrong. Thanks! You're always such a great help. P.S. This is how I drew them XD